OpenStudy (anonymous):

A pair of linear equations is shown below: y = –2x + 3 y = –4x – 1 Which of the following statements best explains the steps to solve the pair of equations graphically?

OpenStudy (anonymous):

@Teddyiswatshecallsme

OpenStudy (anonymous):

@Hell-Scythe

OpenStudy (anonymous):

@TheDream can you help with this one

OpenStudy (anonymous):

What are the statements?

OpenStudy (anonymous):

A. Graph the first equation, which has slope = 3 and y-intercept = –2, graph the second equation, which has slope = –1 and y-intercept = –4, and find the point of intersection of the two lines B. Graph the first equation, which has slope = –3 and y-intercept = 2, graph the second equation, which has slope = 1 and y-intercept = 4 and find the point of intersection of the two lines C. Graph the first equation, which has slope = –2 and y-intercept = 3, graph the second equation, which has slope = –4 and y-intercept = –1, and find the point of intersection of the two lines D. Graph the first equation, which has slope = 2 and y-intercept = –3, graph the second equation, which has slope = 4 and y-intercept = 1, and find the point of intersection of the two lines.

OpenStudy (anonymous):

Equation for a line is y=mx+b where m is slope and b is the y-intercept

OpenStudy (anonymous):

yes I know that

OpenStudy (anonymous):

So which part is tripping you up

OpenStudy (anonymous):

so basically just graph them and explain what it looks like??

OpenStudy (anonymous):

The question says just pick a statement which statement follows the rule y=mx+b

OpenStudy (anonymous):

graph it and see the intersections baby

OpenStudy (anonymous):

Oh ok. Thanks guys

OpenStudy (anonymous):

welcome, baby!

OpenStudy (anonymous):

help me with one more?

OpenStudy (anonymous):

A graph is shown below: Which of the following inequalities is best represented by this graph? A.5x + y ≤ 2 B.5x + y ≥ 2 C.5x – y ≤ 2 D.5x – y ≥ 2

OpenStudy (anonymous):

@.ahamed

OpenStudy (triciaal):

for the 1st one you don't have to graph but you can if you wanted to. the question is just asking for the steps. I think it is just checking if you know the coefficient of the x is the slope and the constant is the intercept. Checking if you know how to do it..

OpenStudy (triciaal):

for the 2nd one one approach is to find the equation first then pick a point and test the solution here y = 5x - 2 same as 2 = 5x - y shaded to the right means > therefore 5x - y >and = 2 example when x = 1 y = 3 y>2